OpenStudy (huangtina):

How to solve this problem. y′=(cos^2x)(cos^2 2y)

OpenStudy (anonymous):

separate the variables and integrate

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\cos^2x*\cos^22y\]

OpenStudy (anonymous):

\[\frac{ 1 }{ \cos^22y }dy = \cos^2x dx\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ \cos^22y }dy = \int\limits_{}^{}\cos^2x dx\]

OpenStudy (priyar):

@huangtina did u understand?

OpenStudy (priyar):

did u get the answer?

OpenStudy (huangtina):

I cannot continue is it right?

OpenStudy (priyar):

your 3rd step from start is wrong.. when u take cos^2 2y to the other side it becomes the reciprocal as wriiten abv..

OpenStudy (priyar):

\[\int\limits_{}^{} \sec^2 (2y) dy = \int\limits_{}^{} \cos^2 x dx\]

OpenStudy (priyar):

got it?

OpenStudy (huangtina):

Thank you for helping!

OpenStudy (priyar):

welcome!